a) A sample of oxygen gas is collected by displacement of water at 25°C and 1.13 atm total pressure. If the vapor pressure of water is 23.756 mm Hg at 25°C, what is the partial pressure of the oxygen gas in the sample?
WebAssign will check your answer for the correct number of significant figures. mm Hg

b) Calcium oxide can be used to "scrub" carbon dioxide from air.
CaO(s) + CO2(g) → CaCO3(s)
What mass of CO2 could be absorbed by 1.38 g of CaO?
WebAssign will check your answer for the correct number of significant figures. g
What volume would this CO2 occupy at STP?
WebAssign will check your answer for the correct number of significant figures. L

c) If water is added to magnesium nitride, ammonia gas is produced when the mixture is heated.
Mg3N2(s) + 3 H2O(l)→ 3 MgO(s) + 2 NH3(g)
If 12.7 g of magnesium nitride is treated with water, what volume of ammonia gas would be collected at 21°C and 739 mm Hg?
WebAssign will check your answer for the correct number of significant figures. L
Answer #1

a) The partial pressure of oxygen in the sample = 1.13*760 mm Hg - 23.756 mm Hg = 835.044 mm Hg

b) Mass of CO2 that could be absorbed by 1.38 g CaO = {1.38 g/(56 g/mol)}*44 g/mol = 1.084 g

c) Volume of ammonia gas collected = nRT/P = (2*12.7/100)*(0.0821 L.atm/mol.K)*(21+273) K/(739/760) atm = 6.305 L